OpenStudy (anonymous):

Lorraine wants to paint the 4 rectangular walls of a room, including the door. The room is 18 feet long and 12 feet wide, and the walls are 10 feet high. What is the total area that will be painted? F 300 ft 2 G 600 ft 2 H 816 ft 2 J 864 ft 2

OpenStudy (anonymous):

my answer not match , Area 4 side?

OpenStudy (anonymous):

18*10*2=360 10*12*2=240 360+240=600 here your answer

OpenStudy (anonymous):

thank you so much